OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 74 and the last term is –100?

OpenStudy (anonymous):

Sn = n/2 [a + l ] = 15 [74 - 100] = -390

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

u welcome